# A Low-Overhead Incorporation-Extrapolation based Few-Shot CSI Feedback Framework for Massive MIMO Systems

### v1

@@ Abstract
Accurate channel state information (CSI) is essential for downlink precoding at the base station (BS), especially for frequency division duplexing (FDD)-based wideband massive multiple-input multiple-output (MIMO) systems with orthogonal frequency-division multiplexing (OFDM). In FDD systems, CSI is attained through CSI feedback from the user equipment (UE). However, large-scale antennas and large number of subcarriers significantly increase CSI feedback overhead. Deep learning-based CSI feedback methods have received tremendous attention in recent years due to their great capability of compressing CSI. Nonetheless, large amounts of collected samples are required to train deep learning models, which is severely challenging in practice. Besides, with the rapidly increasing number of antennas and subcarriers, most of these deep learning methods’ CSI feedback overhead also grow dramatically, owing to their focus on full-dimensional CSI feedback. To address this issue, in this paper, we propose a low-overhead Incorporation-Extrapolation based Few-Shot CSI feedback Framework (IEFSF) for massive MIMO systems. To further reduce the feedback overhead, a low-dimensional eigenvector-based CSI matrix is first formed with the incorporation process at the UE, and then recovered to the full-dimensional eigenvector-based CSI matrix at the BS via the extrapolation process. After that, to alleviate the necessity of the extensive collected samples and enable few-shot CSI feedback, we further propose a knowledge-driven data augmentation (KDDA) method and an artificial intelligence-generated content (AIGC) -based data augmentation method by exploiting the domain knowledge of wireless channels and by exploiting a novel generative model, respectively. The KDDA method is applicable for scenarios with very few collected samples (e.g., only 100 - 200 samples) while the AIGC-based method exhibits better performance when there are several hundred samples. Experimental results are also provided based on the DeepMIMO dataset. Numerical results demonstrate that the proposed IEFSF can significantly reduce CSI feedback overhead by $16$ times compared with existing CSI feedback methods while maintaining higher feedback accuracy using only several hundreds of collected samples.
@@ I Introduction
Although there are vector quantization-based and codebook-based CSI feedback methods for conventional FDD MIMO systems, both of them face significant challenges in FDD massive MIMO systems[7], especially under the wideband bandwidth configuration. As the number of antennas at the BS in massive MIMO systems increasingly increases and the operational bandwidth progressively widens, the dimension of the CSI matrix grows significantly, leading to unaffordable high feedback overhead. To reduce the feedback overhead, in recent years, numerous deep learning-based CSI feedback methods have been proposed to compress and feedback the CSI matrix at the UE and then reconstruct it at the BS, including CsiNet [8], CsiNet+[9], SRNet[10], TransNet [11], etc. CsiNet, known as the first deep learning-based CSI feedback method, was proposed to compress and recover CSI by learning the channel structure from training samples with convolutional neural networks (CNNs). In [9], CsiNet+ was developed from CsiNet by considering multiple-rate bit-level compressive sensing for CSI feedback. SRNet in [10] employed the image super-resolution technology to exploit CSI frequency characteristics. In [11], TransNet, a neural network for CSI feedback based on the Transformer architecture [12], was presented. TransNet can learn connections between angular domain elements of the CSI and thus showed significant performance improvement compared to the existing deep learning-based CSI feedback methods. Note that, to efficiently reduce the CSI feedback overhead, all these aforementioned deep learning-based CSI feedback methods are typically designed with an encoder-decoder architecture, i.e., the UE-side encoder compresses the CSI matrix and then feedbacks the compressed CSI to BS through the feedback link, and the BS-side decoder reconstructs the CSI matrix after receiving the compressed one. In addition, the CSI feedback efficiency can be further promoted by exploiting the channel property. For example, in [13], the SampleDL method was proposed via capitalizing on channel correlations in both temporal and frequency domains. Based on the channel correlations, only a portion of the CSI elements was selected and a much size-reduced massive MIMO CSI matrix was obtained before the CSI feedback. This, therefore, effectively reduced the feedback overhead and computational complexity on the UE side. Nonetheless, such downsampling process will inevitably lead to information loss due to the direct dropping of the orthogonal frequency-division multiplexing (OFDM) symbols and the subcarriers.
@@
Furthermore, previous research on deep learning-based CSI feedback has primarily focused on full channel feedback, which involves compressing the full downlink channel matrix on the UE side and subsequently reconstructing it on the BS side after feedback. In contrast, as specified in current 3rd Generation Partnership Project (3GPP) specification[14], the commonly used codebook-based CSI feedback is based on an implicit CSI feedback scheme. That is, only partial channel information (e.g., the precoding matrix indicator (PMI)) is encoded with an encoding module at the UE, fed back to the BS, and decoded with a decoding module at the BS. In view of the increasing antenna scale and transmission bandwidth, CSI feedback based on PMI will be greatly helpful to reduce the feedback overhead. As a result, eigenvector-based CSI feedback methods were proposed in [15, 16, 17], where eigenvectors were used as the PMI for downlink precoding.
@@
Despite the effectiveness of these deep learning-based CSI feedback methods, a large number of training samples (e.g., tens of thousands of samples) are generally required during the training period to obtain deep learning models with strong generalization ability. For example, in [8], $100,000$ collected samples were required to train the CsiNet under the system configuration of $32$ antennas and $1024$ subcarriers. In practice, collecting such large amount of samples necessitates labor-intensive and time-consuming channel measurement campaigns, which is always unaffordable. In addition, the mobility of UEs poses another challenge to data collection due to the fast-changing channel conditions and synchronization issues. To circumvent such data collection problem, learning from a merely limited number of samples is highly advocated, which is known as few-shot learning (FSL) in machine learning. To realize FSL, effectively augmenting the limited collected samples is an essential approach. Existing data augmentation methods that have been used for CSI feedback can be roughly categorized into three categories, i.e., the image processing-based, the knowledge-driven, and the artificial intelligence-generated content (AIGC) -based data augmentation methods. The image processing-based methods regard each CSI matrix from the CSI samples as an image with two channels (typically the real and the imaginary parts of the complex-valued CSI matrix). Therefore, traditional image data augmentation methods, e.g., scaling, shifting, and rotation, can be applied to CSI samples directly[18]. However, without proper guidance and evaluation, such image processing-based data augmentation methods may destroy the original structural information of CSI samples (e.g., the orientation of the antenna array), which may introduce nuisances contaminating the model training. The knowledge-driven data augmentation methods aim to explore domain knowledge of wireless channels to generate augmented samples. For example, in [19], two knowledge-driven data augmentation methods were proposed by considering the geographical continuity of antennas and the delay property of MIMO channels. Recently, with the emergence of ChatGPT[20], Stable Diffusion[21], DALL-E-2[22], etc., AIGC, which indicates content generated with generative models, has been garnering increasing attention from both academia and industry [23, 24]. Exploiting AIGC for data augmentation becomes promising in various fields[25, 26, 27]. Specifically, generative adversarial networks (GANs), one kind of generative model for AIGC, have been studied for data augmentation in wireless communications. GANs first model wireless channels by learning the distribution from a set of channel realizations[28, 29, 30]. After channel modeling, the trained GANs could generate various channel realizations following the same distribution to enhance the performance and generalization ability of deep learning-based CSI feedback methods[30]. Nevertheless, training a GAN typically requires large amounts of training data[31], which contradicts the intention of exploiting GAN for data augmentation, especially under few-shot scenarios. Therefore, a practical AIGC-based data augmentation method for CSI feedback deserves further exploration.
@@
    We propose the incorporation-extrapolation CSI feedback scheme to simultaneously reduce feedback overhead and computational complexity. Unlike [13], which directly downsamples the frequency domain subcarriers, the proposed incorporation-extrapolation scheme first yields a low-dimensional eigenvector-based CSI matrix through the incorporation process by dividing all subcarriers into several groups and generating eigenvectors for all groups. Then, a Transformer-based CSI feedback network is proposed for CSI compression and reconstruction by exploiting the channel correlation among multiple subcarrier groups. Next, a novel lightweight frequency extrapolation network is proposed to recover the full-dimensional eigenvector-based CSI matrix on the BS side with low computational complexity by exploiting multi-path extrapolation and residual learning. [^1] Note that the ‘multi-path’ extrapolation indicates extrapolation through multiple extrapolation modules, which differs from the concept of ‘multipath’ propagation phenomenon in wireless signal propagation.
@@
    We evaluate the performance of the proposed IEFSF under both indoor and outdoor scenarios. Numerical results indicate that both the proposed knowledge-driven data augmentation method and the proposed AIGC-based data augmentation method exhibit excellent performance in achieving accurate CSI feedback with only several hundreds of collected samples under the system configuration of $32$ antennas and $100$ MHz bandwidth with $1024$ subcarriers. In addition, the proposed IEFSF can achieve a higher CSI feedback accuracy than existing CSI feedback methods but with $16$ times less feedback overhead.
@@ II-A System Model
To maximize system performance, the downlink precoding matrix $\mathbf{P}_{i}$ needs to be designed based on the downlink channel $\mathbf{H}_{i}$. However, in FDD systems, the BS cannot directly derive $\mathbf{H}_{i}$ from the uplink channel information obtained through uplink pilot training. In addition, the increasing number of antennas and subcarriers lead to a significant growth of the channel dimension, making the overhead of full channel feedback become prohibitive. As a result, implicit feedback mechanism (e.g., PMI feedback) is advocated by current 3GPP specifications. In this paper, we consider eigenvalue decomposition (EVD) -based downlink precoding and thus the eigenvector-based CSI feedback is adopted as in [15, 16, 17]. Specifically, denote the Gram matrix of $\mathbf{H}_{i}$ as $\mathbf{R}_{i}\in\mathbb{C}^{N_{T}\times N_{T}}$, i.e.,
@@
Then, the eigenvector-based CSI matrix $\mathbf{W}$, instead of the whole downlink channel $\mathbf{H}\in\mathbb{C}^{N_{R}\times(N_{T}\times N_{c})}\triangleq[\mathbf{H}_{1},\mathbf{H}_{2},\ldots,\mathbf{H}_{N_{c}}]$, will be fed back to the BS, and the $i$-th column vector of the reconstructed eigenvector-based CSI matrix $\hat{\mathbf{W}}$ will serve as the downlink precoding vector $\mathbf{p}_{i}$ at the $i$-th subcarrier.
@@ II-B CSI Feedback Procedure
In most of existing deep learning-based CSI feedback methods, after obtaining $\mathbf{W}$, the UE starts the CSI feedback process by first compressing $\mathbf{W}$ into the CSI feature vector $\mathbf{v}\in\mathbb{R}^{L_{q}\times 1}$ via the compression module, then quantizing $\mathbf{v}$ to the bit stream via the quantization module, and finally sending the bit stream to the BS via the feedback link. Once receiving the bit stream, the BS first dequantizes the bit stream to the CSI feature vector $\hat{\mathbf{v}}\in\mathbb{R}^{L_{q}\times 1}$ with the dequantization module, [^4] Note that quantization refers to using a limited number of bits to represent each element of $\mathbf{v}$ with loss of precision, and dequantization refers to the inverse operation of quantization to recover $\mathbf{v}$ from the acquired bit stream[32, 33]. and then reconstructs the eigenvector-based CSI matrix with the reconstruction module.
@@ III-A Incorporation Process
Thanks to the frequency domain correlation among adjacent subcarriers, we can derive a low-dimensional CSI matrix from $\mathbf{W}$ for CSI feedback by incorporating the information of neighboring subcarriers together. Note that, although there is correlation between adjacent subcarriers, the specific channel information (e.g., angle of departure, channel gain, etc) at different subcarriers may still differ from each other[34], especially when the subcarrier spacing is large. Therefore, different from SampleDL which directly drops subcarriers, our proposed incorporation process is designed to merge the channel information of adjacent subcarriers together to try to preserve the channel information as much as possible. Specifically, denoting the set containing the channel matrices of all subcarriers as $\mathcal{H}_{subc}=\{\mathbf{H}_{i}|i=1,2,\ldots,N_{c}\}$, we first divide all subcarriers into $N_{grp}=N_{c}/N_{gr}$ groups, i.e., $\mathcal{H}_{grp}^{j}$, $j=1,2,\ldots,N_{grp}$, with a given granularity $N_{gr}$. [^5] For simplicity, we assume that $N_{c}$ is divisible by $N_{gr}$. Concretely, the $j$-th group $\mathcal{H}_{grp}^{j}$ can be expressed by
@@ III-B CSI Compression and Reconstruction
where $F_{\text{en}}$ represents the encoder composed of the compression and quantization modules and parameterized by $\mathbf{\Theta}_{\text{de}}=\mathbf{\Theta}_{\text{com}}\cup\mathbf{\Theta}_{\text{quan}}$, and $F_{\text{de}}$ represents the decoder composed of the dequantization and reconstruction modules and parameterized by $\mathbf{\Theta}_{\text{de}}=\mathbf{\Theta}_{\text{rec}}\cup\mathbf{\Theta}_{\text{deq}}$.
Figure 1: The incorporation-extrapolation CSI feedback scheme which consists of the incorporation process, the proposed Transformer-based CSI feedback network, and the frequency extrapolation network.
Note that with an appropriate $N_{gr}$, the frequency domain correlation can be retained at the group level. Therefore, by exploiting the group-level frequency domain correlation, the CSI matrix at adjacent subcarrier groups can be effectively compressed and reconstructed. In the proposed framework, we leverage the Transformer network to explore the correlation among subcarrier groups for CSI feedback. As shown in Fig. 1, the compression module and the reconstruction module of the Transformer-based CSI feedback network include $N_{com}$ and $N_{rec}$ identical layers, respectively. Each of these identical layers consists of two components: a multi-head self-attention (MHSA) sub-layer and a feed-forward (FF) sub-layer. The MHSA sub-layer performs computations on different subcarrier groups to generate a self-attention matrix. This self-attention matrix plays a crucial role in paying more attention to correlated inputs, thereby improving the effectiveness of the high-dimensional feature learning of the CSI matrix. Such improved feature learning facilitates the subsequent CSI compression and reconstruction operations. The FF sub-layer is designed to enhance the non-linear representation capability of the layer. To further optimize the training process, residual connection (RC)[35] and layer normalization (LN)[36] operations are applied after both the MHSA sub-layer and the FF sub-layer. These operations help improve the neural network’s training efficiency. Additionally, since the MHSA sub-layer treats all subcarrier groups equally and does not consider their positional information, supplementary positional encodings (PEs) are inserted before the compression module and the reconstruction module to compensate for this limitation, such that the positional information can also be utilized for better CSI compression and reconstruction. Finally, we exploit uniform quantization (UQ) and corresponding uniform dequantization (UDQ) for conversion between the bit stream and CSI feature vector in the proposed Transformer-based CSI feedback network.

### v2

@@ Abstract
Accurate channel state information (CSI) is essential for downlink precoding in frequency division duplexing (FDD) massive multiple-input multiple-output (MIMO) systems with orthogonal frequency-division multiplexing (OFDM). However, obtaining CSI through feedback from the user equipment (UE) becomes challenging with the increasing scale of antennas and subcarriers and leads to extremely high CSI feedback overhead. Deep learning-based methods have emerged for compressing CSI but these methods generally require substantial collected samples and thus pose practical challenges. Moreover, existing deep learning methods also suffer from dramatically growing feedback overhead owing to their focus on full-dimensional CSI feedback. To address these issues, we propose a low-overhead Incorporation-Extrapolation based Few-Shot CSI feedback Framework (IEFSF) for massive MIMO systems. An incorporation-extrapolation scheme for eigenvector-based CSI feedback is proposed to reduce the feedback overhead. Then, to alleviate the necessity of extensive collected samples and enable few-shot CSI feedback, we further propose a knowledge-driven data augmentation (KDDA) method and an artificial intelligence-generated content (AIGC) -based data augmentation method by exploiting the domain knowledge of wireless channels and by exploiting a novel generative model, respectively. Experimental results based on the DeepMIMO dataset demonstrate that the proposed IEFSF significantly reduces CSI feedback overhead by 64 times compared with existing methods while maintaining higher feedback accuracy using only several hundred collected samples.
@@ I Introduction
Although there are conventional vector quantization-based and codebook-based CSI feedback methods for FDD MIMO systems, both of them face significant challenges in FDD massive MIMO systems[7], especially under the wideband bandwidth configuration. As the number of antennas at the BS in massive MIMO systems increasingly increases and the operational bandwidth progressively widens, the dimension of the CSI matrix grows significantly, leading to unaffordable high feedback overhead. To reduce the feedback overhead, in recent years, numerous deep learning-based CSI feedback methods have been proposed to compress and feedback the CSI matrix at the UE and then reconstruct it at the BS, including CsiNet [8], CsiNet+[9], SRNet[10], TransNet [11], etc. CsiNet, known as the first deep learning-based CSI feedback method, was proposed to compress and recover CSI by learning the channel structure from training samples with convolutional neural networks (CNNs). In [9], CsiNet+ was developed from CsiNet by considering multiple-rate bit-level compressive sensing for CSI feedback. SRNet in [10] employed the image super-resolution technology to exploit CSI frequency characteristics. In [11], TransNet, a neural network for CSI feedback based on the Transformer architecture [12], was presented. TransNet can learn connections between angular domain elements of the CSI and thus showed significant performance improvement compared to the existing deep learning-based CSI feedback methods. Note that, to efficiently reduce the CSI feedback overhead, all these aforementioned deep learning-based CSI feedback methods are typically designed with an encoder-decoder architecture, i.e., the UE-side encoder compresses the CSI matrix and then feedbacks the compressed CSI to BS through the feedback link, and the BS-side decoder reconstructs the CSI matrix after receiving the compressed one. In addition, the CSI feedback efficiency can be further promoted by exploiting the channel property. For example, in [13], the SampleDL method was proposed via capitalizing on channel correlations in both temporal and frequency domains. Based on the channel correlations, only a portion of the CSI elements was selected and a much size-reduced massive MIMO CSI matrix was obtained before the CSI feedback. This, therefore, effectively reduced the feedback overhead and computational complexity on the UE side. Nonetheless, such downsampling process will inevitably lead to information loss due to the direct dropping of the orthogonal frequency-division multiplexing (OFDM) symbols and the subcarriers.
@@
Furthermore, previous research on deep learning-based CSI feedback has primarily focused on full channel feedback, which involves compressing the full downlink channel matrix on the UE side and subsequently reconstructing it on the BS side after feedback. In contrast, as specified in the current 3rd Generation Partnership Project (3GPP) specification[14], the commonly used codebook-based CSI feedback is based on an implicit CSI feedback scheme. That is, only partial channel information (e.g., the precoding matrix indicator (PMI)) is encoded with an encoding module at the UE, fed back to the BS, and decoded with a decoding module at the BS. In view of the increasing antenna scale and transmission bandwidth, CSI feedback based on PMI will be greatly helpful in reducing the feedback overhead. As a result, eigenvector-based CSI feedback methods were proposed in [15, 16, 17], where eigenvectors were used as the PMI for downlink precoding.
@@
Despite the effectiveness of these deep learning-based CSI feedback methods, a large number of collected samples (e.g., tens of thousands of samples) are generally required during the training period to obtain deep learning models with strong generalization ability. For example, in [8], $100,000$ collected samples were required to train the CsiNet under the system configuration of $32$ antennas and $1024$ subcarriers. In practice, collecting such large amounts of samples necessitates substantial CSI acquisition time and pilot training overhead [18], which is always unaffordable. In addition, confined by the generalization capability of existing DL-based CSI feedback methods, it is also necessary for these methods to collect new data for new scenarios, making the extensive data collection impractical for employment. Moreover, UE mobility poses another challenge to data collection, i.e., the fast-changing channel conditions originating from the mobility necessitate much more time-frequency resources to obtain accurate CSI samples. To circumvent such data collection problems, learning from a merely limited number of collected samples is highly advocated, which is known as few-shot learning (FSL) in machine learning. To realize FSL, effectively augmenting the limited collected samples is an essential approach. Existing data augmentation methods that have been used for CSI feedback can be roughly categorized into four categories, i.e., the image processing-based, the ray-tracing based, the knowledge-driven, the artificial intelligence-generated content (AIGC) -based data augmentation methods. The image processing-based methods regard each CSI matrix from the CSI samples as an image with two channels (typically the real and the imaginary parts of the complex-valued CSI matrix). Therefore, traditional image data augmentation methods, e.g., scaling, shifting, and rotation, can be applied to CSI samples directly[19]. However, without proper guidance and evaluation, such image processing-based data augmentation methods may destroy the original structural information of CSI samples (e.g., the orientation of the antenna array), which may introduce nuisances contaminating the model training. Recently, ray-tracing based data augmentation methods have shown the potential to augment CSI samples by utilizing digital twin techniques[20, 21]. However, ray-tracing based data augmentation methods essentially require a digital reconstruction of exact wireless communication environments to make the generated CSI samples similar to the measured CSI samples (i.e., following the same distribution), which is significantly challenging and computationally intensive in practice. Inaccurate digital reconstruction might produce dissimilar samples and ultimately degrade the model performance. In particular, for mobile scenarios, it is even impossible to accurately model the mobility of mobile users, making ray-tracing based data augmentation methods less beneficial. The knowledge-driven data augmentation methods aim to explore domain knowledge of wireless channels to generate augmented samples. For example, in [22], two knowledge-driven data augmentation methods were proposed by considering the geographical continuity of antennas and the delay property of MIMO channels. Recently, with the emergence of ChatGPT[23], Stable Diffusion[24], DALL-E-2[25], etc., AIGC, which indicates content generated with generative models, has been garnering increasing attention from both academia and industry [26, 27]. Exploiting AIGC for data augmentation becomes promising in various fields[28, 29, 30]. Specifically, generative adversarial networks (GANs), one kind of generative model for AIGC, have been studied for data augmentation in wireless communications. GANs first model wireless channels by learning the distribution from a set of channel realizations[31, 32, 33]. After channel modeling, the trained GANs could generate various channel realizations following the same distribution to enhance the performance and generalization ability of deep learning-based CSI feedback methods[33]. Nevertheless, training a GAN typically requires large amounts of training data[34], which contradicts the intention of exploiting GAN for data augmentation, especially under few-shot scenarios. Therefore, a practical AIGC-based data augmentation method for CSI feedback deserves further exploration.
@@
    We propose the incorporation-extrapolation CSI feedback scheme to simultaneously reduce feedback overhead and computational complexity. Unlike [13], which directly downsamples the frequency domain subcarriers, the proposed incorporation-extrapolation scheme first yields a low-dimensional eigenvector-based CSI matrix through the incorporation process by dividing all subcarriers into several groups and generating eigenvectors for all groups. Then, a Transformer-based CSI compression and reconstruction network is proposed for CSI compression and reconstruction by exploiting the channel correlation among multiple subcarrier groups. Next, a novel lightweight frequency extrapolation network is proposed to recover the full-dimensional eigenvector-based CSI matrix on the BS side with low computational complexity by exploiting multi-path extrapolation and residual learning. [^1] Note that the ‘multi-path’ extrapolation indicates extrapolation through multiple extrapolation modules, which differs from the concept of ‘multipath’ propagation phenomenon in wireless signal propagation. The proposed extrapolation network significantly reduces the computational complexity of frequency domain extrapolation compared with existing extrapolation networks.
@@
    We evaluate the performance of the proposed IEFSF under both indoor and outdoor scenarios. Numerical results indicate that both the proposed knowledge-driven data augmentation method and the proposed AIGC-based data augmentation method exhibit excellent performance in achieving accurate CSI feedback with only several hundreds of collected samples under the system configuration of $32$ antennas and $100$ MHz bandwidth with $1024$ subcarriers. In addition, the proposed IEFSF can achieve a higher CSI feedback accuracy than existing CSI feedback methods but with $64$ times less feedback overhead.
@@ II-A System Model
To maximize system performance, the downlink precoding matrix $\mathbf{P}_{i}$ needs to be designed based on the downlink channel $\mathbf{H}_{i}$. However, in FDD systems, the BS cannot directly derive $\mathbf{H}_{i}$ from the uplink channel information obtained through uplink pilot training. In addition, the increasing number of antennas and subcarriers leads to a significant growth of the channel dimension, making the overhead of full channel feedback become prohibitive. As a result, an implicit feedback mechanism (e.g., PMI feedback) is advocated by current 3GPP specifications. In this paper, we consider eigenvalue decomposition (EVD) -based downlink precoding and thus the eigenvector-based CSI feedback is adopted as in [15, 16, 17]. Specifically, denote the Gram matrix of $\mathbf{H}_{i}$ as $\mathbf{R}_{i}\in\mathbb{C}^{N_{T}\times N_{T}}$, i.e.,
@@
Then, the eigenvector-based CSI matrix $\mathbf{W}$, instead of the whole downlink channel $\mathbf{H}\triangleq[\mathbf{H}_{1},\mathbf{H}_{2},\ldots,\mathbf{H}_{N_{c}}]\in\mathbb{C}^{N_{R}\times(N_{T}\times N_{c})}$, will be fed back to the BS, and the $i$-th column vector of the reconstructed eigenvector-based CSI matrix $\hat{\mathbf{W}}$ will serve as the downlink precoding vector $\mathbf{p}_{i}$ at the $i$-th subcarrier.
@@ II-B CSI Feedback Procedure
In most existing deep learning-based CSI feedback methods, after obtaining $\mathbf{W}$, the UE starts the CSI feedback process by first compressing $\mathbf{W}$ into the CSI feature vector $\mathbf{v}\in\mathbb{R}^{L_{q}\times 1}$ via the compression module, then quantizing $\mathbf{v}$ to the bit stream via the quantization module, and finally sending the bit stream to the BS via the feedback link. Once receiving the bit stream, the BS first dequantizes the bit stream to the CSI feature vector $\hat{\mathbf{v}}\in\mathbb{R}^{L_{q}\times 1}$ with the dequantization module, [^4] Note that quantization refers to using a limited number of bits to represent each element of $\mathbf{v}$ with loss of precision, and dequantization refers to the inverse operation of quantization to recover $\mathbf{v}$ from the acquired bit stream[35, 36]. and then reconstructs the eigenvector-based CSI matrix with the reconstruction module.
@@ III-A Incorporation Process
Thanks to the frequency domain correlation among adjacent subcarriers, we can derive a low-dimensional CSI matrix from $\mathbf{W}$ for CSI feedback by incorporating the information of neighboring subcarriers together. Note that, although there is a correlation between adjacent subcarriers, the specific channel information (e.g., angle of departure, channel gain, etc) at different subcarriers may still differ from each other[37], especially when the subcarrier spacing is large. Therefore, different from SampleDL which directly drops subcarriers, our proposed incorporation process is designed to merge the channel information of adjacent subcarriers together to try to preserve the channel information as much as possible. Specifically, denoting the set containing the channel matrices of all subcarriers as $\mathcal{H}_{subc}=\{\mathbf{H}_{i}|i=1,2,\ldots,N_{c}\}$, we first divide all subcarriers into $N_{grp}=N_{c}/N_{gr}$ groups, i.e., $\mathcal{H}_{grp}^{j}$, $j=1,2,\ldots,N_{grp}$, with a given granularity $N_{gr}$. [^5] For simplicity, we assume that $N_{c}$ is divisible by $N_{gr}$. Concretely, the $j$-th group $\mathcal{H}_{grp}^{j}$ can be expressed by
@@ III-B CSI Compression and Reconstruction
where $F_{\text{en}}$ represents the encoder composed of the compression and quantization modules and parameterized by $\mathbf{\Theta}_{\text{en}}=\mathbf{\Theta}_{\text{com}}\cup\mathbf{\Theta}_{\text{quan}}$, and $F_{\text{de}}$ represents the decoder composed of the dequantization and reconstruction modules and parameterized by $\mathbf{\Theta}_{\text{de}}=\mathbf{\Theta}_{\text{rec}}\cup\mathbf{\Theta}_{\text{deq}}$.
Figure 1: The incorporation-extrapolation CSI feedback scheme which consists of the incorporation process, the proposed Transformer-based CSI compression and reconstruction network, and the frequency extrapolation network.
Note that with an appropriate $N_{gr}$, the frequency domain correlation can be retained at the group level. Therefore, by exploiting the group-level frequency domain correlation, the CSI matrix at adjacent subcarrier groups can be effectively compressed and reconstructed. In the proposed framework, we leverage the Transformer network to explore the correlation among subcarrier groups for CSI feedback. As shown in Fig. 1, the compression module and the reconstruction module of the Transformer-based CSI compression and reconstruction network include $N_{com}$ and $N_{rec}$ identical layers, respectively. Each of these identical layers consists of two components: a multi-head self-attention (MHSA) sub-layer and a feed-forward (FF) sub-layer. The MHSA sub-layer performs computations on different subcarrier groups to generate a self-attention matrix. This self-attention matrix plays a crucial role in paying more attention to correlated inputs, thereby improving the effectiveness of the high-dimensional feature learning of the CSI matrix. Such improved feature learning facilitates the subsequent CSI compression and reconstruction operations. The FF sub-layer is designed to enhance the non-linear representation capability of the layer. To further optimize the training process, residual connection (RC)[38] and layer normalization (LN)[39] operations are applied after both the MHSA sub-layer and the FF sub-layer. These operations help improve the neural network’s training efficiency. Additionally, since the MHSA sub-layer treats all subcarrier groups equally and does not consider their positional information, supplementary positional encodings (PEs) are inserted before the compression module and the reconstruction module to compensate for this limitation, such that the positional information can also be utilized for better CSI compression and reconstruction. Finally, we exploit uniform quantization (UQ) and corresponding uniform dequantization (UDQ) for conversion between the bit stream and CSI feature vector in the proposed Transformer-based CSI compression and reconstruction network.